OpenStudy (anonymous):

S = { (1,2,3), (4,5,6), (11,16,21) } Does the vector (6,9,12) \[\in\] span{S}

OpenStudy (anonymous):

I solved the matrix: x y z b [ 1 4 11 | 6 ] [ 2 5 16 | 9 ] [ 3 6 21 | 12]

ganeshie8 (ganeshie8):

You can eyeball it

OpenStudy (anonymous):

and got x=2-3z, y=1-2z and z in R

OpenStudy (anonymous):

What do those answers tell me?

ganeshie8 (ganeshie8):

that tells you that you can reach the vector (6, 9, 12) by taking linear combinations of the given vectors

ganeshie8 (ganeshie8):

-(1,2,3)-(4,5,6)+(11,16,21) = (6,9,12)

OpenStudy (anonymous):

The moment you said to eyeball it, i immediately noticed that.

ganeshie8 (ganeshie8):

they cook up these examples precicely for eyeballing

OpenStudy (anonymous):

Next time i'll take a look before solving like crazy, thanks ganeshie8

ganeshie8 (ganeshie8):

np :)